1.find the point of intersection between the lines:<3,-1,2>+t<1,1,-1> and <-8,2,0>+t<-3,2,-7>.

2.show that the lines x+1=3t,y=1,z+5=2t for tER and z+2=s,y-3=-5s,z+4=-2s for tER intersect, and find the point of intersection.

3.find the point of intersection between the planes: -5x+y-2z=3 and 2x-3y+5z=-7.


1
Expert's answer
2021-06-30T06:40:52-0400

1.


"\begin{matrix}
 x=3+t,& y=-1+t,& z=2-t \\
 x=-8-3s, &y=2+2s, & z=-7s
\end{matrix}"


"\begin{matrix}
 3+t=-8-3s \\
 -1+t=2+2s \\
2-t=-7s
\end{matrix}"

"\begin{matrix}
 3+t=-8-3s \\
 4=-10-5s \\
1=2-5s
\end{matrix}"


"\begin{matrix}
 3+t=-8-3s \\
 5s=-14\\
5s=1
\end{matrix}"

No solution. Therefore the given lines do not intersect.


2.


"\begin{matrix}
 x=-1+3t,& y=1,& z=-5+2t \\
 x=-2+s, &y=3-5s, & z=-4-2s
\end{matrix}"


"\begin{matrix}
 -1+3t=-2+s \\
 1=3-5s \\
-5+2t=-4-2s
\end{matrix}"

"\begin{matrix}
 s=0.4 \\
 t=-0.2\\
t=0.1
\end{matrix}"

"\begin{matrix}
 3+t=-8-3s \\
 5s=-14\\
5s=1
\end{matrix}"

No solution. Therefore the given lines do not intersect.


3.


"\begin{matrix}
 -5x+y-2z=3\\
2x-3y+5z=-7
\end{matrix}"


"\begin{matrix}
 y=5x+2z+3 \\
 -13x-z=2
\end{matrix}"

"\begin{matrix}
 y=-21x-1 \\
 z=-13x-2
\end{matrix}"

We can write our parametrization of the line of intersection as


"\langle0,-1,-2\rangle+t\langle1,-21, -13\rangle, t\in \R"